Theobromine is a bitter alkaloid of cacao plant. It is found in chocolate, leaves of tea plant and cola nut. How many grams of theobromine (C7H8N4O2) should be dissolved in 375g water to make 3.75 m theobromine solution?


1
Expert's answer
2021-03-04T06:37:30-0500

Solution:


C7H8N4O2 = 180.164 "(\frac{g}{mole})"


(3.75 "\frac{mole}{kg}" )*(0.375 kg)*(180.164 "(\frac{g}{mole})" ) = 253 g of C7H8N4O2